OpenStudy (anonymous):

abcd is a rectangle. find the length of each diagonal. ac=2(5a+1) bd=2(a+1)

OpenStudy (anonymous):

if abcd is a rectangle the diagonals are the same length

OpenStudy (anonymous):

what is the answer though

OpenStudy (anonymous):

it could be either - are u sure you have all the info on this question? i assume these are alternatives?

OpenStudy (anonymous):

thats all the question says

OpenStudy (anonymous):

sorry i cant help u -

OpenStudy (anonymous):

diagonal is 4.5 unit do u need working???

OpenStudy (anonymous):

how did you find it and can you help me with one more? same question but different numbers ac=3y/5 bd=3y-4

OpenStudy (anonymous):

Hey Ing, I will give detailed explanation for the first one. That will help you to do the rest on your own. Is that ok???

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

Hey Ing, i thnik u had typed the first question wrong so i will use second question for explanation

OpenStudy (anonymous):

if u draw rectangle ABCD, you will find that its diagonals are AC AND BD In a rectangle the diagonals are ALWAYS EQUAL, so AC=BD this means 3y ---- = 3y-4 5 3y = 5(3y-4) 3y = 15y -20 or 15y-20=3y 15y-3y=20 12y=20 y=20/12 y=5/3

OpenStudy (anonymous):

now putting y=5/3 in ac and bd, we get 3 x 5/3 15/3 5 ac = --------- = -------- = ------ = 1 5 5 5 5 15 bd = 3y-4 = 3 x ---- - 4 = ------ - 4 = 5-4 = 1 3 3

OpenStudy (anonymous):

Is it clear Ing ?????? If yes, pls click the GOOD ANSWER button :}